Lewis Companys standard labor cost of producing one unit of Product DD is 3.4 hours at the rate of $13.8 per hour. During August, 43,700 hours of labor are incurred at a cost of $14.00 per hour to produce 12,600 units of Product DD Compute the total labor variance Total labor variance Compute the labor price and quantity variances. Labor price variance Labor quantity variance Compute the labor price and quantity variances, assuming the standard is 3.8 hours of direct labor at $14.20 per hour. Labor price variance Labor quantity variance
Answer #1

a. Total Labour variance = (AH * AR) - (SH * SR) =  (43700 * 14) - (42840# * 13.80) = $20,608 U

# 12600 * 3.40

b. Labour Price Variance = (AH * AR) - (AH * SR) = (43700 * 14) - (43700 * 13.80) = $8,740 U

Labour Quantity Variance = (AH * SR) - (SH * SR) = (43,700 * 13.80) - (42840# * 13.80) = $11,868 U

# 12600 * 3.40

c. Labour Price Variance = (AH * AR) - (AH * SR) = (43700 * 14) - (43700 * 14.20) = $8,740 F

Labour Quantity Variance = (AH * SR) - (SH * SR) = (43,700 * 14.20) - (47880# * 14.20) = $59,356 F

# 12600 * 3.80